# A Covariant and Manifestly Projective Invariant Formulation of Thomas-Whitehead Gravity

### v1

@@ Appendix A Conventions and Definitions
$\hat{\mathcal{K}}_{d}^{\ abc}=$ $\mathcal{K}^{\bar{d}}_{\ \bar{a}\bar{b}\bar{c}}\mathcal{G}_{d\bar{d}}^{\ \ \ a\bar{a}[b|\bar{b}|c]\bar{c}}$ (A.44)
$\mathcal{G}_{d\bar{d}}^{\ \ \ a\bar{a}b\bar{b}c\bar{c}}=$ $G_{d\bar{d}}g^{a\bar{a}}g^{b\bar{b}}g^{c\bar{c}}-4\delta^{b}_{\ d}\delta^{\bar{b}}_{\ \bar{d}}g^{a\bar{a}}g^{c\bar{c}}+\delta^{b}_{\ d}\delta^{\bar{b}}_{\ \bar{d}}g^{ac}g^{\bar{a}\bar{c}}$ (A.45)
$\breve{\mathcal{K}}^{abc}=$ $g^{ap}g^{bm}g^{cn}\breve{\nabla}_{[m}{\mathcal{D}}_{n]p}~{}~{}~{}.$ (A.46)
Appendix B Recasting the Original TW Field Equations in a Manifestly Tensorial Form
@@
$\tfrac{1}{\sqrt{|g|}}\breve{\nabla}_{a}(\sqrt{|g|}Anything)=$ (˘$(\breve{\nabla}_{a}-({\rm d}+1)g_{a})Anything~{}~{}~{}.$ (B.1)
B.1 Recasting $\mathbb{D}^{ab}$ in a Manifestly Tensorial Form
Equation (B.1) allows us to expand terms such as
$\tfrac{1}{\sqrt{|g|}}\breve{\nabla}_{a}(\sqrt{|g|}K^{(ab)c})=$ (˘$(\breve{\nabla}_{a}-({\rm d}+1)g_{a})K^{(ab)c}$ (B.2)
$=$ $(\nabla_{a}-({\rm d}+2)\text{\textcent}_{c})K^{(ab)c}+g_{c}K^{(ab)c}$ (B.3)
where we have used Eqs. (A.2) and (A.7) to produce the last line. It will also be of use to expand Eqs. (A.44) and (A.46)
$\hat{{\mathcal{K}}}_{c}{}^{(ab)d}=$ $2{\mathcal{K}}_{c}{}^{(ab)d}-2\lambda_{0}^{2}g_{c}g_{m}{\mathcal{K}}^{m(ab)d}+\delta^{d}{}_{c}(8{\mathcal{K}}^{ab}-2g^{ab}{\mathcal{K}})+\delta^{(a}{}_{c}(g^{b)d}{\mathcal{K}}-4{\mathcal{K}}^{b)d})$ (B.4)
$\breve{{\mathcal{K}}}^{(ab)c}=$ $K^{(ab)c}-g_{m}{\mathcal{K}}^{m(ab)c}~{}~{}~{}.$ (B.5)
In expanding $\breve{{\mathcal{K}}}^{(ab)c}$, we have made use of the definition of the projective Cotton-York tensor from Eq. (2.40). Contracting indices in Eq. B.4 and simplifying yields the following useful result
$\tfrac{1}{2}\hat{{\mathcal{K}}}_{c}{}^{(ab)c}=-\lambda_{0}^{2}g_{c}g_{m}{\mathcal{K}}^{m(ab)c}+2(2{\rm d}-3){\mathcal{K}}^{ab}-({\rm d}-1)g^{ab}{\mathcal{K}}~{}~{}~{}.$ (B.6)
Recall the field equation $\mathbb{D}^{ab}$ in its manifestly projectively invariant form
$\mathbb{D}^{ab}\equiv-\tfrac{({\rm d}-1)}{4\kappa_{0}J_{0}}g^{ab}+\tfrac{1}{2}\hat{\mathcal{K}}_{c}^{\ (ab)c}+\tfrac{\lambda_{0}^{2}}{\sqrt{|g|}}\breve{\nabla}_{c}(\sqrt{|g|}K^{(ab)c})-\lambda_{0}^{2}g_{c}\breve{\mathcal{K}}^{(ab)c}=0~{}~{}~{}.$ (2.59)
Utilizing the expansions in Eqs. (B.2), (B.6), and (B.5) and simplifying casts Eq. $\mathbb{D}^{ab}$ into its manifestly tensorial form
$\mathbb{D}^{ab}=$ $-\tfrac{d-1}{4J_{0}\kappa_{0}}g^{ab}-(d-1)g^{ab}\mathcal{K}+2(2d-3)\mathcal{K}^{ab}+\lambda_{0}^{2}(\nabla_{c}-({\rm d}+2)\text{\textcent}_{c})K^{(ab)c}=0.$ (2.70)
Utilizing Eq. (3.4), it is straightforward to show the last term collapses to an overall ¯$\bar{\nabla}_{c}$ derivative. This reduces the equation to the covariant and manifestly projectively invariant form
$\mathbb{D}^{ab}=$ $-\tfrac{d-1}{4J_{0}\kappa_{0}}g^{ab}-(d-1)g^{ab}\mathcal{K}+2(2d-3)\mathcal{K}^{ab}+\lambda_{0}^{2}\bar{\nabla}_{c}K^{(ab)c}~{}~{}~{}.$ (3.40)
B.2 Recasting $\mathbb{H}^{ab}$ in a Manifestly Tensorial Form
Using Eqs. (B.1), (A.2), and (A.7) we find the following identity that we will find useful in this section
$\tfrac{1}{\sqrt{|g|}}\breve{\nabla}_{a}(\sqrt{|g|}{\mathcal{K}}^{a}{}_{bcd}K^{bcd})=$ $(\nabla_{a}-({\rm d}+1)\text{\textcent}_{a}){\mathcal{K}}^{a}{}_{bcd}K^{bcd}~{}~{}~{}.$ (B.7)
@@ B.2 Recasting $\mathbb{H}^{ab}$ in a Manifestly Tensorial Form
$\mathcal{H}_{mn}=$ $-\mathcal{K}_{mn}-\kappa_{0}g_{mn}\left[\tfrac{2J_{0}\lambda_{0}^{2}}{({\rm d}+1)\sqrt{|g|}}\breve{\nabla}_{a}(\sqrt{|g|}{\mathcal{K}}^{a}{}_{bcd}K^{bcd})+\mathcal{L}\right]$ (B.8)
$+2\kappa_{0}J_{0}\lambda_{0}^{2}(K_{mcd}K_{n}{}^{cd}+2K_{bcm}K^{bc}{}_{n})+2\kappa_{0}J_{0}(8{\mathcal{K}}_{mb}{\mathcal{K}}^{b}{}_{n}-2{\mathcal{K}}{\mathcal{K}}_{mn})$ (B.9)
$+2\kappa_{0}J_{0}({\mathcal{K}}_{mbcd}{\mathcal{K}}_{n}{}^{bcd}-{\mathcal{K}}^{a}{}_{mcd}{\mathcal{K}}_{an}{}^{cd}-2{\mathcal{K}}^{abc}{}_{m}{\mathcal{K}}_{abcn})$ (B.10)
[…]
$\mathcal{L}=$ $-\tfrac{1}{2\kappa_{0}}{\mathcal{K}}+J_{0}\lambda_{0}^{2}K_{bcd}K^{bcd}-J_{0}(\mathcal{K}_{abcd}\mathcal{K}^{abcd}-4\mathcal{K}_{bd}\mathcal{K}^{bd}+\mathcal{K}^{2})$ (2.69)
we use Eq. (B.7) along with Eqs. (A), (A), (A), and (A.39) and simplify to cast $\mathcal{H}_{mn}$ into the following form
$\hat{R}_{mn}-\tfrac{1}{2}g_{mn}\hat{R}=\kappa_{0}\hat{T}_{mn}$ (B.11)
$\hat{T}_{mn}=-\frac{1}{\kappa_{0}}\Big{(}Q_{mn}-\tfrac{1}{2}g_{mn}Q+({\rm d}-1)\mathcal{P}_{mn}-{\rm d}{\mathcal{P}}_{[mn]}-\tfrac{{\rm d}-1}{2}g_{mn}\mathcal{P}\Big{)}$
$\qquad-J_{0}g_{mn}\left[\lambda_{0}^{2}K_{bcd}K^{bcd}-({\mathcal{K}}_{abcd}{\mathcal{K}}^{abcd}-4{\mathcal{K}}_{ab}{\mathcal{K}}^{ab}+{\mathcal{K}}^{2})\right]$ (B.12)
$\qquad+g_{mn}2J_{0}\lambda_{0}^{2}(-\tfrac{1}{{\rm d}+1}\nabla_{a}+\text{\textcent}_{a})(K_{bcd}\mathcal{K}^{abcd})$ (B.13)
$\qquad+2J_{0}\lambda_{0}^{2}(K_{mcd}K_{n}{}^{cd}+2K_{bcm}K^{bc}{}_{n})+2J_{0}(8{\mathcal{K}}_{mb}{\mathcal{K}}^{b}{}_{n}-2{\mathcal{K}}{\mathcal{K}}_{mn})$ (B.14)
$\qquad+2J_{0}({\mathcal{K}}_{mbcd}{\mathcal{K}}_{n}{}^{bcd}-{\mathcal{K}}^{a}{}_{mcd}{\mathcal{K}}_{an}{}^{cd}-2{\mathcal{K}}^{abc}{}_{m}{\mathcal{K}}_{abcn})]~{}~{}~{}.$
@@
Combining lines one and two of $\hat{T}_{mn}$ and simplifying results in the covariant form of Einstein’s equations
$\hat{R}_{mn}-\tfrac{1}{2}g_{mn}\hat{R}=\kappa_{0}\hat{T}_{mn}$ (2.71)
$\hat{T}_{mn}=-\frac{1}{\kappa_{0}}\Big{(}Q_{mn}-\tfrac{1}{2}g_{mn}Q+({\rm d}-1)\mathcal{P}_{mn}-{\rm d}{\mathcal{P}}_{[mn]}-\tfrac{{\rm d}-1}{2}g_{mn}\mathcal{P}\Big{)}$
$\qquad+J_{0}g_{mn}\left[2\lambda_{0}^{2}(-\tfrac{1}{{\rm d}+1}\nabla_{a}+\text{\textcent}_{a})K_{bcd}\mathcal{K}^{abcd}-\lambda_{0}^{2}K^{bcd}K_{bcd}\right.$ (B.15)
$\left.\qquad\qquad\qquad+\mathcal{K}_{abcd}\mathcal{K}^{abcd}-4\mathcal{K}^{ab}\mathcal{K}_{ab}+\mathcal{K}^{2}\right]+2J_{0}\lambda_{0}^{2}(K_{mcd}K_{n}{}^{cd}+2K_{bcm}K^{bc}{}_{n})$
$\qquad+2J_{0}(8{\mathcal{K}}_{mb}{\mathcal{K}}^{b}{}_{n}-2{\mathcal{K}}{\mathcal{K}}_{mn})+2J_{0}({\mathcal{K}}_{mbcd}{\mathcal{K}}_{n}{}^{bcd}-{\mathcal{K}}^{a}{}_{mcd}{\mathcal{K}}_{an}{}^{cd}-2{\mathcal{K}}^{abc}{}_{m}{\mathcal{K}}_{abcn}).$
@@
B.3 Recasting $\tilde{E}_{c}{}^{ab}$ in terms of Tensorial Objects to Demonstrate it is not a Tensor
@@
Using Eqs. (B.1) and (B.2) it is straightforward to prove the following identities that we will find useful in this section
$\tfrac{1}{\sqrt{|g|}}\breve{\nabla}_{c}(\sqrt{|g|}T^{ab})=$ $(\nabla_{c}-2\text{\textcent}_{c})T^{ab}-\delta^{(a}{}_{c}\text{\textcent}_{m}T^{b)m}-({\rm d}-1)g_{c}T^{ab}+\delta^{(a}{}_{c}g_{m}T^{b)m}$ (B.16)
$\tfrac{1}{\sqrt{|g|}}\breve{\nabla}_{d}(\sqrt{|g|}K_{c}{}^{(ab)d})=$ $(\nabla_{d}-({\rm d}+1)\text{\textcent}_{d}){\mathcal{K}}_{c}{}^{(ab)d}-2\text{\textcent}_{c}{\mathcal{K}}^{ab}+2g_{c}{\mathcal{K}}^{ab}$ (B.17)
$\tfrac{1}{\sqrt{|g|}}\breve{\nabla}_{d}(\sqrt{|g|}g_{c}K^{(ab)d})=$ $K^{(ab)d}(\breve{\nabla}_{c}g_{d}+g_{c}g_{d})+g_{c}(\nabla_{d}-({\rm d}+2)\text{\textcent}_{d})K^{(ab)d}~{}~{}~{}.$ (B.18)
In the above, $T^{ab}=T^{ba}$ represents an arbitrary symmetric tensor. Starting with the original form of the equations of motion
~$\tilde{\mathcal{E}}_{c}{}^{ab}\equiv\mathcal{E}_{c}^{\ ab}-\tfrac{1}{d+1}\delta_{c}^{\ (a}\mathcal{E}_{d}^{\ b)d}=0$ (2.63)
$\mathcal{E}_{c}^{\ ab}=\mathcal{E}_{c}^{\ ba}\equiv\tfrac{1}{4\kappa_{0}J_{0}\sqrt{|g|}}\breve{\nabla}_{c}(\sqrt{|g|}g^{ab})-\tfrac{1}{2\sqrt{|g|}}\breve{\nabla}_{d}(\sqrt{|g|}\hat{\mathcal{K}}_{c}^{\ (ab)d})\quad$
$\qquad\qquad\qquad\qquad+\tfrac{\lambda_{0}^{2}}{\sqrt{|g|}}\breve{\nabla}_{d}(\sqrt{|g|}g_{c}\breve{\mathcal{K}}^{(ab)d})-\lambda_{0}^{2}K^{(ab)d}\mathcal{D}_{dc}~{}~{}~{}$ (2.4)
@@
we use Eqs. (B.4) and (B.5) to substitute for $\hat{{\mathcal{K}}}_{c}{}^{(ab)d}$ and $\breve{{\mathcal{K}}}^{(ab)d}$. After some simplification this results in
$\mathcal{E}_{c}^{\ ab}=$ $\tfrac{1}{\sqrt{|g|}}\breve{\nabla}_{c}\left[\sqrt{|g|}\left((\tfrac{1}{4J_{0}\kappa_{0}}+\mathcal{K})g^{ab}-4\mathcal{K}^{ab}\right)\right]-\tfrac{1}{\sqrt{|g|}}\breve{\nabla}_{d}(\sqrt{|g|}\mathcal{K}_{c}^{\ (ab)d})\quad$ (B.19)
@@
$+\lambda_{0}^{2}\left[\tfrac{1}{\sqrt{|g|}}\breve{\nabla}_{d}(\sqrt{|g|}g_{c}K^{(ab)d})-{\mathcal{D}}_{cd}K^{(ab)d}\right]$ (B.20)
@@
$-\tfrac{1}{2}\delta^{(a}{}_{c}\tfrac{1}{\sqrt{|g|}}\breve{\nabla}_{d}\left[\sqrt{|g|}\left(g^{b)d}{\mathcal{K}}-4{\mathcal{K}}^{b)d}\right)\right]~{}~{}~{}.$ (B.21)
@@
Next, we substitute using Eqs. (B.16), (B.17), and (B.18) and simplify to the final form
$\mathcal{E}_{c}{}^{ab}=$ $\mathbb{E}_{c}{}^{ab}+g_{c}\mathbb{D}^{ab}+\delta^{(a}{}_{c}\mathfrak{g}^{b)}$ (2.73)
$\mathbb{E}_{c}{}^{ab}\equiv$ $(\nabla_{c}-2\text{\textcent}_{c})\mathscr{K}^{ab}-\delta^{(a}{}_{c}\text{\textcent}_{d}\mathscr{K}^{b)d}-[(\nabla_{d}-({\rm d}+1)\text{\textcent}_{d})\mathcal{K}_{c}{}^{(ab)d}-2\text{\textcent}_{c}\mathcal{K}^{ab}]$ (B.22)
$-\lambda_{0}^{2}K^{(ab)d}(\mathcal{P}_{cd}-\nabla_{c}\text{\textcent}_{d}-\text{\textcent}_{c}\text{\textcent}_{d})$ (2.74)
$\mathscr{K}^{ab}\equiv$ $(\tfrac{1}{4J_{0}\kappa_{0}}+\mathcal{K})g^{ab}-4\mathcal{K}^{ab}$ (2.76)
@@
$\mathfrak{g}^{b}\equiv$ $\tfrac{1}{4J_{0}\kappa_{0}}g_{d}g^{bd}-\tfrac{1}{2}(\nabla_{d}-({\rm d}+3)\text{\textcent}_{d})(g^{bd}\mathcal{K}-4\mathcal{K}^{bd})$ (2.77)
@@
where in the last line of the definition of $\mathbb{E}_{c}{}^{ab}$, we have made use of Eq. (A.42) to switch between ${\mathcal{D}}_{cd}$ and ${\mathcal{P}}_{cd}$. Subtracting the trace then results in the full equation of motion
~$\tilde{\mathcal{E}}_{c}{}^{ab}=$ $\mathbb{E}_{c}{}^{ab}-\tfrac{1}{{\rm d}+1}\delta^{(a}{}_{c}\mathbb{E}_{m}{}^{b)m}+g_{c}\mathbb{D}^{ab}-\tfrac{1}{{\rm d}+1}\delta^{(a}{}_{c}g_{m}\mathbb{D}^{b)m}=0~{}~{}~{}.$ (2.78)

### v3

@@ Appendix A Conventions and Definitions
$\hat{\mathcal{K}}_{d}^{\ abc}=\mathcal{K}^{\bar{d}}_{\ \bar{a}\bar{b}\bar{c}}\mathcal{G}_{d\bar{d}}^{\ \ a\bar{a}[b|\bar{b}|c]\bar{c}}$ (308)
$\mathcal{G}_{d\bar{d}}^{\ \ \ a\bar{a}b\bar{b}c\bar{c}}$
$\hskip 3.0pt=G_{d\bar{d}}g^{a\bar{a}}g^{b\bar{b}}g^{c\bar{c}}-4\delta^{b}_{\ d}\delta^{\bar{b}}_{\ \bar{d}}g^{a\bar{a}}g^{c\bar{c}}+\delta^{b}_{\ d}\delta^{\bar{b}}_{\ \bar{d}}g^{ac}g^{\bar{a}\bar{c}}$ (309)
$\breve{\mathcal{K}}^{abc}=g^{ap}g^{bm}g^{cn}\breve{\nabla}_{[m}{\mathcal{D}}_{n]p}~{}.$ (310)
Note that the object $\hat{{\mathcal{K}}}_{d}^{\ abc}$ simply has $G_{d\bar{d}}$ in its definition as on $\mathcal{M}$ indices are raised or lowered via $g^{ma}$ or $g_{ma}$ as
${\mathcal{K}}_{abcd}\equiv$ $g_{ma}{\mathcal{K}}^{m}{}_{bcd}$ (311)
${\mathcal{K}}^{abcd}=$ $g^{bp}g^{cm}g^{dn}{\mathcal{K}}^{a}{}_{pmn}$ (312)
$K^{bcd}\equiv$ $g^{bp}g^{cm}g^{dn}K_{pmn}~{}.$ (313)
Important relationships between the projective curvatures and the Levi-Civita curvatures are
$\mathcal{K}^{a}{}_{bcd}=$ $\hat{R}^{a}_{\ bcd}+\tilde{Q}^{a}_{\ bcd}+\delta^{a}{}_{[c}\bar{\mathcal{P}}_{d]b}$ (314)
$=$ $\hat{R}^{a}_{\ bcd}+Q^{a}_{\ bcd}+\delta^{a}{}_{[c}\mathcal{P}_{d]b}-\delta^{a}{}_{b}\mathcal{P}_{[cd]}$ (315)
$\mathcal{K}_{mn}=$ $\hat{R}_{mn}+\tilde{Q}_{mn}+({\rm d}-1)\bar{\mathcal{P}}_{mn}$ (316)
$=$ $\hat{R}_{mn}+Q_{mn}+({\rm d}-1)\mathcal{P}_{mn}-{\rm d}\mathcal{P}_{[mn]}$ (317)
$\mathcal{K}=$ $\hat{R}+\tilde{Q}+({\rm d}-1)\bar{\mathcal{P}}$ (318)
$=$ $\hat{R}+Q+({\rm d}-1)\mathcal{P}~{}.$ (319)
These allow for straightforward transitions to the form of Einstein equations as well as transitions between formulations.
Appendix B Recasting the Original TW Field Equations in a Covariant Form
@@
$\tfrac{1}{\sqrt{|g|}}\breve{\nabla}_{a}(\sqrt{|g|}~{}\text{anything})=(\breve{\nabla}_{a}-({\rm d}+1)g_{a})\text{anything.}$. (320)
B.1 Recasting $\mathbb{D}^{ab}$ in a Covariant Form
Equation (320) allows us to expand terms such as
$\tfrac{1}{\sqrt{|g|}}\breve{\nabla}_{c}(\sqrt{|g|}K^{(ab)c})=(\breve{\nabla}_{c}-({\rm d}+1)g_{c})K^{(ab)c}$ (321)
$\qquad\qquad\qquad=(\nabla_{c}-({\rm d}+2)\text{\textcent}_{c})K^{(ab)c}+g_{c}K^{(ab)c}~{}~{}~{}~{}~{}~{}~{}$ (322)
where we have used Eqs. (262) and (267) to produce the last line. It will also be of use to expand Eqs. (308) and (310)
$\hat{{\mathcal{K}}}_{c}^{\ (ab)d}=2{\mathcal{K}}_{c}^{\ (ab)d}-2\lambda_{0}^{2}g_{c}g_{m}{\mathcal{K}}^{m(ab)d}$
$\qquad+\delta^{d}_{\ c}(8{\mathcal{K}}^{ab}-2g^{ab}{\mathcal{K}})+\delta^{(a}_{\ \ c}(g^{b)d}{\mathcal{K}}-4{\mathcal{K}}^{b)d})$ (323)
$\breve{{\mathcal{K}}}^{(ab)c}=K^{(ab)c}-g_{m}{\mathcal{K}}^{m(ab)c}~{}.$ (324)
Note that $\hat{{\mathcal{K}}}_{d}^{\ abc}$, that includes $G_{d\bar{d}}$ in its definition, should not be confused with ${\mathcal{K}}_{d}^{\ abc}=g_{dm}{\mathcal{K}}^{mabc}$ that, here on $\mathcal{M}$, has its indices raised or lowered via $g^{dm}$ or $g_{dm}$ as in Eq. (312). In expanding $\breve{{\mathcal{K}}}^{(ab)c}$ above, we have made use of the definition of the projective Cotton-York tensor from Eq. (46).
Contracting indices in Eq. B.1 and simplifying yields the following useful result
$\tfrac{1}{2}\hat{{\mathcal{K}}}_{c}{}^{(ab)c}=$ $-\lambda_{0}^{2}g_{c}g_{m}{\mathcal{K}}^{m(ab)c}+2(2{\rm d}-3){\mathcal{K}}^{ab}$ (325)
$\qquad\qquad\qquad\qquad-({\rm d}-1)g^{ab}{\mathcal{K}}~{}.$ (326)
Recall the field equation $\mathbb{D}^{ab}$ in its manifestly projective invariant form
$0=\mathbb{D}^{ab}\equiv$ $-\tfrac{({\rm d}-1)}{4J_{0}\kappa_{0}}g^{ab}+\tfrac{1}{2}\hat{\mathcal{K}}_{c}^{\ (ab)c}$ (327)
$+\tfrac{\lambda_{0}^{2}}{\sqrt{|g|}}\breve{\nabla}_{c}(\sqrt{|g|}K^{(ab)c})-\lambda_{0}^{2}g_{c}\breve{\mathcal{K}}^{(ab)c}.$ (69)
Utilizing the expansions in Eqs. (321), (325), and (324) and simplifying casts Eq. $\mathbb{D}^{ab}$ into its covariant form
$0=\mathbb{D}^{ab}=\lambda_{0}^{2}(\nabla_{c}-({\rm d}+2)\text{\textcent}_{c})K^{(ab)c}-({\rm d}-1)\mathbb{K}^{ab}-2\mathcal{K}^{ab}$ (81)
$\hskip 6.0pt\mathbb{K}^{ab}\equiv g^{ab}\left({\mathcal{K}}+\tfrac{1}{4J_{0}\kappa_{0}}\right)-4\mathcal{K}^{ab}~{}.$ (328)
Via use of Eq. (106), it is straightforward to show the first term collapses to an overall ¯$\bar{\nabla}_{c}$ derivative. This reduces the equation to the covariant and manifestly projective invariant form
$0=\mathbb{D}^{ab}=$ $\lambda_{0}^{2}\bar{\nabla}_{c}K^{(ab)c}-({\rm d}-1)\mathbb{K}^{ab}-2\mathcal{K}^{ab}~{}.$ (158)
B.2 Recasting $\mathcal{H}_{mn}$ in a Covariant Form
Using Eqs. (320), (262), and (267) we find the following identity that we will find useful in this section
$\tfrac{1}{\sqrt{|g|}}\breve{\nabla}_{a}(\sqrt{|g|}{\mathcal{K}}^{a}_{\ bcd}K^{bcd})=(\nabla_{a}-({\rm d}+1)\text{\textcent}_{a})({\mathcal{K}}^{a}_{\ bcd}K^{bcd}).$ (329)
@@ B.2 Recasting $\mathcal{H}_{mn}$ in a Covariant Form
$0=\mathcal{H}_{mn}\equiv$ $-\mathcal{K}_{mn}+\kappa_{0}\left(2\mathbb{L}_{mn}-g_{mn}\mathcal{L}\right)$ (330)
$-g_{mn}\tfrac{2J_{0}\kappa_{0}\lambda_{0}^{2}}{({\rm d}+1)\sqrt{|g|}}\breve{\nabla}_{a}(\sqrt{|g|}{\mathcal{K}}^{a}{}_{bcd}K^{bcd})$ (76)
$\mathcal{L}=$ $-\frac{1}{2\kappa_{0}}(\mathcal{K}+2\Lambda_{0})+\mathbb{L}$ (54)
@@
we use Eq. (329) along with Eqs. (316) and (318). Simplifying yields the covariant form of Einstein’s equations
$\hat{R}_{mn}-$ $\tfrac{1}{2}g_{mn}\hat{R}=\kappa_{0}\hat{T}_{mn}+g_{mn}\Lambda_{0}$ (83)
$\hat{T}_{mn}=$ $-\kappa_{0}^{-1}\left(Q_{mn}-\tfrac{1}{2}g_{mn}Q\right)+2\mathbb{L}_{mn}-g_{mn}\mathbb{L}$ (331)
$-\kappa_{0}^{-1}\left(({\rm d}-1)\mathcal{P}_{mn}-{\rm d}{\mathcal{P}}_{[mn]}-\tfrac{{\rm d}-1}{2}g_{mn}\mathcal{P}\right)$ (332)
$-g_{mn}\tfrac{2J_{0}\lambda_{0}^{2}}{{\rm d}+1}(\nabla_{a}-({\rm d}+1)\text{\textcent}_{a})(K_{bcd}\mathcal{K}^{abcd})~{}.$ (84)
@@
B.3 Recasting ~$\tilde{\mathcal{E}}_{c}{}^{ab}$ in terms of Tensorial Objects to Demonstrate it is not a Tensor
@@
Using Eqs. (320) and (321) it is straightforward to prove the following identities that we will find useful in this section
$\tfrac{1}{\sqrt{|g|}}\breve{\nabla}_{c}(\sqrt{|g|}T^{ab})=$ $(\nabla_{c}-2\text{\textcent}_{c})T^{ab}-\delta^{(a}_{\ \ c}\text{\textcent}_{m}T^{b)m}$ (333)
$-({\rm d}-1)g_{c}T^{ab}+\delta^{(a}_{\ \ c}g_{m}T^{b)m}$ (334)
$\tfrac{1}{\sqrt{|g|}}\breve{\nabla}_{d}(\sqrt{|g|}K_{c}{}^{(ab)d})=$ $(\nabla_{d}-({\rm d}+1)\text{\textcent}_{d}){\mathcal{K}}_{c}{}^{(ab)d}$ (335)
$-2\text{\textcent}_{c}{\mathcal{K}}^{ab}+2g_{c}{\mathcal{K}}^{ab}$ (336)
$\tfrac{1}{\sqrt{|g|}}\breve{\nabla}_{d}(\sqrt{|g|}g_{c}K^{(ab)d})=$ $K^{(ab)d}(\breve{\nabla}_{c}g_{d}+g_{c}g_{d})$ (337)
$+g_{c}(\nabla_{d}-({\rm d}+2)\text{\textcent}_{d})K^{(ab)d}.$ (338)
In the above, $T^{ab}=T^{ba}$ represents an arbitrary symmetric tensor.
Starting with the original form of the equations of motion
$0=\tilde{\mathcal{E}}_{c}{}^{ab}\equiv$ $\mathcal{E}_{c}^{\ ab}-\tfrac{1}{d+1}\delta_{c}^{\ (a}\mathcal{E}_{d}^{\ b)d}$ (74)
$\mathcal{E}_{c}^{\ ab}\equiv$ $\tfrac{1}{4J_{0}\kappa_{0}\sqrt{|g|}}\breve{\nabla}_{c}(\sqrt{|g|}g^{ab})-\tfrac{1}{2\sqrt{|g|}}\breve{\nabla}_{d}(\sqrt{|g|}\hat{\mathcal{K}}_{c}^{\ (ab)d})$
$+\tfrac{\lambda_{0}^{2}}{\sqrt{|g|}}\breve{\nabla}_{d}(\sqrt{|g|}g_{c}\breve{\mathcal{K}}^{(ab)d})-\lambda_{0}^{2}K^{(ab)d}\mathcal{D}_{dc}$ (II.4)
@@
we use Eqs. (B.1) and (324) to substitute for $\hat{{\mathcal{K}}}_{c}{}^{(ab)d}$ and $\breve{{\mathcal{K}}}^{(ab)d}$. After some simplification this results in
$\mathcal{E}_{c}^{\ ab}=$ $\tfrac{1}{\sqrt{|g|}}\breve{\nabla}_{c}(\sqrt{|g|}\mathbb{K}^{ab})-\tfrac{1}{\sqrt{|g|}}\breve{\nabla}_{d}(\sqrt{|g|}\mathcal{K}_{c}^{\ (ab)d})\quad$
@@
$+\lambda_{0}^{2}\left[\tfrac{1}{\sqrt{|g|}}\breve{\nabla}_{d}(\sqrt{|g|}g_{c}K^{(ab)d})-{\mathcal{D}}_{cd}K^{(ab)d}\right]$
@@
$-\tfrac{1}{2}\delta^{(a}{}_{c}\tfrac{1}{\sqrt{|g|}}\breve{\nabla}_{d}\left[\sqrt{|g|}\left(g^{b)d}{\mathcal{K}}-4{\mathcal{K}}^{b)d}\right)\right]$ (339)
@@
where $\mathbb{K}^{ab}$ is as in Eq. (328). Next, we substitute using Eqs. (333), (335), and (337) and simplify to the final form
$\mathcal{E}_{c}{}^{ab}=$ $\mathbb{E}_{c}{}^{ab}+g_{c}\mathbb{D}^{ab}+\delta^{(a}{}_{c}\mathfrak{g}^{b)}$ (87)
$\mathbb{E}_{c}{}^{ab}\equiv$ $(\nabla_{c}-2\text{\textcent}_{c})\mathbb{K}^{ab}-\delta^{(a}{}_{c}\text{\textcent}_{d}\mathbb{K}^{b)d}$ (340)
$-[(\nabla_{d}-({\rm d}+1)\text{\textcent}_{d})\mathcal{K}_{c}{}^{(ab)d}-2\text{\textcent}_{c}\mathcal{K}^{ab}]$ (341)
$-\lambda_{0}^{2}(\mathcal{P}_{cd}-\nabla_{c}\text{\textcent}_{d}-\text{\textcent}_{c}\text{\textcent}_{d})K^{(ab)d}$ (88)
@@
$\mathfrak{g}^{b}\equiv$ $\tfrac{1}{4J_{0}\kappa_{0}}g_{d}g^{bd}-\tfrac{1}{2}(\nabla_{d}-({\rm d}+3)\text{\textcent}_{d})(g^{bd}\mathcal{K}-4\mathcal{K}^{bd}).~{}~{}~{}~{}~{}~{}$ (91)
@@
where in the last line of the definition of $\mathbb{E}_{c}{}^{ab}$, we have made use of Eq. (304) to switch between ${\mathcal{D}}_{cd}$ and ${\mathcal{P}}_{cd}$. Subtracting the trace results in a form of the equation that is clearly non-tensorial
$0=\tilde{\mathcal{E}}_{c}^{\ ab}=$ $\tilde{\mathbb{E}}_{c}{}^{ab}+g_{c}\mathbb{D}^{ab}-\tfrac{1}{{\rm d}+1}\delta^{(a}_{\ \ c}g_{m}\mathbb{D}^{b)m}$ (92)
$\tilde{\mathbb{E}}_{c}{}^{ab}\equiv$ $\mathbb{E}_{c}{}^{ab}-\tfrac{1}{{\rm d}+1}\delta_{c}{}^{(a}\mathbb{E}_{m}{}^{b)m}~{}.$ (93)